# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v4

@@ 4.2 Geographical Distribution of Research
Figure 3: Geographical distribution of published papers.
Figure 3 visualizes the geographical distribution of research contributions. China leads with 164 publications, followed by the United States (74) and Australia (32). This dominance underscores the strong focus of Chinese institutions on AI-driven SCRA, supported by extensive investments in AI research and SC optimization. The global spread of contributions emphasizes the universal relevance of SCRA, with increasing participation from emerging economies.
@@
Gupta S 5
Kumar A 4
Modgil S 4
Bag S 3
Hu XJ 3
Wang GJ 3
Xie C 3
Yao G 3
Zhu Y 3
Antoniou G 2
@@ 4.3 Analysis of Authors
As shown in Table  2, the top 10 authors in the field have significantly influenced the domain of SCRA using AI/ML. “Gupta S” leads with 10 publications, followed by “Kumar A” (8) and “Liu YK” (7). These authors represent thought leaders who have contributed extensively to developing AI-based risk assessment frameworks and methodologies.
@@ 4.4 Publisher Contributions
Figure 4: Publisher-wise distribution of the reviewed articles.
Figure 4 presents the contributions of various publishers. Elsevier is the most prominent, contributing over 120 articles, followed by Springer Nature and MDPI. These publishers serve as key platforms for disseminating high-quality research, reflecting their alignment with the evolving needs of AI and SCRA scholarship.
4.5 Journal Contributions
Figure 5: Journal-wise distribution of the reviewed articles.
In Figure 5, journal contributions are analyzed, with “Sustainability” being the most active source of publications, followed by “International Journal of Production Research” and “Annals of Operations Research.” These journals provide vital platforms for researchers to share AI-driven SC risk management advancements.
@@ 4.6 Co-citation Analysis
Co-citation analysis was conducted to explore the intellectual structure of AI-SCRA research, focusing on authors, references, and publication sources. Figure 6(a) illustrates a network of 55 authors grouped into three clusters, emphasizing the pivotal roles of researchers such as Ivanov and Baryannis, who have significantly shaped the development of AI-driven risk assessment frameworks. Figure 6(b) focuses on cited references, identifying five distinct clusters highlighting recurring themes, including adopting hybrid AI models, developing advanced risk prediction frameworks, and exploring post-COVID SCR strategies. Lastly, Figure 6(c) underscores the centrality of academic journals, with the International Journal of Production Research standing out as a key publication venue, reflecting its substantial impact on the field’s scholarly discourse and research progression.
@@ 4.7 Co-authorship Analysis
Co-authorship analysis is a powerful method for understanding collaborations among authors and researchers based on joint publication activities. This analysis highlights the collaborative nature of the research domain, uncovering influential partnerships and the significant contributions of specific countries.
Figure 7(a) illustrates the co-authorship network among authors in this field. Using a threshold of at least one document and a minimum of 55 citations per author, 48 authors out of a total of 304 were identified as significant contributors. The visualization, generated using VOSViewer, reveals four distinct clusters, color-coded as red, green, blue, and yellow. These clusters represent groups of authors with strong collaborative ties. For example, the robust connections surrounding “Modgil” in the red cluster indicate frequent co-authorship and shared research initiatives, reflecting a high degree of collaboration within the cluster.
Figure 7(b) showcases the co-authorship network at the country level. By applying thresholds of at least two documents and nine citations per country, 18 out of 36 countries were included in the analysis. The visualization highlights England and China as leading contributors, dominating in terms of document count, citation impact, and total link strength. This dominance underscores their pivotal roles in advancing AI/ML-driven SCRA research. The results also emphasize strong international collaborations between these countries and other key players, fostering a globally interconnected research ecosystem.
@@ 4.8 Co-occurrence Analysis
The co-occurrence analysis offers valuable insights into the thematic structure of SCRA and AI/ML research by identifying how keywords are interconnected within the scholarly literature. Using VOSviewer, we conducted this analysis by focusing on the units of analysis, such as “Authors’ Keywords” and “Keywords Plus.” This approach reveals latent connections and thematic clusters, enabling a deeper understanding of this domain’s critical topics and prevalent themes.
The network map shown in Figure 8(a) illustrates the relationships among the 26 most frequently occurring author keywords. These keywords were selected from an initial pool of 1,483, applying a minimum threshold of three occurrences. Larger nodes, such as “artificial intelligence” and “machine learning,” indicate higher frequencies. The proximity of the nodes reflects their degree of association, with closely grouped nodes forming thematic clusters. Four clusters emerged, highlighting key themes, including risk management, sustainability, and digital transformation.
Figure 8(b) illustrates the co-occurrence of keywords found in the reviewed literature. The analysis identified 47 frequently occurring terms from a total of 1,032, using a minimum threshold of three occurrences. These keywords were categorized into five distinct clusters, reflecting prevalent themes and connections within the research field. Terms such as “management” and “framework” exhibited the strongest linkages, indicating their crucial roles in connecting different thematic areas. Additionally, other significant clusters highlighted emerging trends, including “supply chain resilience,” “blockchain,” and “sustainability.” This co-occurrence analysis offers a comprehensive overview of the intellectual structure of SCRA research, showcasing both established and emerging themes. The findings emphasize the interconnectivity of concepts and provide guidance for future research directions by identifying critical areas of scholarly focus.
@@
{tblr}
column2 = c, column3 = c, hline1-2,13 = -, Affiliations & Published Paper Count % of 51 articles
NEOMA BUSINESS SCHOOL 5 5.319
CHINESE ACADEMY OF SCIENCES 4 4.255
EMLYON BUSINESS SCHOOL 4 4.255
HEFEI UNIVERSITY OF TECHNOLOGY 4 4.255
HONG KONG POLYTECHNIC UNIVERSITY 4 4.255
HUNAN UNIVERSITY 4 4.255
INTERNATIONAL MANAGEMENT INSTITUTE (IMI) KOLKATA 4 4.255
UNIVERSITY OF CAMBRIDGE 4 4.255
UNIVERSITY OF TECHNOLOGY SYDNEY 4 4.255
UNIVERSITY OF HUDDERSFIELD 3 3.191
OTHERS 214 57.45
@@ 4.9 Top 10 Affiliations
Table  3 highlights the top 10 affiliations contributing to SCRA research, ranked by their published paper count. These institutions collectively demonstrate significant academic leadership, with NEOMA Business School leading with 5 publications (5.319% of the total reviewed articles). Several institutions, including the Chinese Academy of Sciences, Emlyon Business School, and Hong Kong Polytechnic University, each contributed 4 papers representing 4.255% of the reviewed literature. These affiliations underscore a global research effort, with prominent contributions spanning Europe, Asia, and Australia. Together, the top affiliations account for a substantial share of foundational studies in SCRA, promoting advancements in the application of AI to mitigate SC risks. Meanwhile, the category labeled “Others” collectively represents the remaining 57.45% of contributions, reflecting a broad and diverse research landscape.
@@ 4.10.1 Authors
Figure 9 illustrates the bibliographic coupling analysis of 28 authors, grouping them into six thematic clusters based on shared references, highlighting intellectual connections in SCR research. Prominent contributors like Ivanov D, Gupta S, and Wang G-J demonstrate centrality, reflecting their significant influence in areas such as risk prediction, disruption management, and quantitative modeling. The clusters reveal interdisciplinary themes, including resilience strategies, industry-specific applications, and the integration of AI-driven methods. This analysis showcases both established research directions and emerging trends, providing a roadmap for future studies and collaborations in SCR.
@@ 4.10.2 Countries
Figure 10 illustrates the bibliographic coupling among 18 countries, categorizing them into clusters highlighting global research collaborations in AI-driven SCRA. Countries such as China, the United States, and Germany emerge as central hubs, showcasing their pivotal roles in advancing research and promoting international partnerships. The clusters signify thematic alignments across regions, with nations collaborating on critical topics like predictive analytics, sustainability, and resilience strategies. This visualization underscores the significance of cross-border collaboration in addressing global SC challenges through AI innovation.
@@ 4.10.3 Documents
Figure 11 illustrates the bibliographic coupling analysis of 62 documents, visually depicting the thematic clustering and citation relationships among the studies. The eight clusters highlight distinct research themes within the field, showcasing the interconnectedness of key works. Highly influential documents, such as those by “Baryannis G.” and “Ivanov D.,” emerge as central nodes, reflecting their significant contributions to the discourse on SCRA using AI. This visualization underscores the collaborative and evolving nature of the research domain, providing insights into how various studies align and contribute to specific thematic advancements.
@@ 4.10.4 Organizations
Figure 12 presents the bibliographic coupling analysis of 24 organizations, highlighting their collaborative and thematic alignments in AI-based SCRA research. The three clusters represent distinct institutional networks, with prominent contributors like the Hong Kong Polytechnic University and the Indian Institute of Technology Delhi playing central roles. This visualization emphasizes the global distribution of research efforts and showcases the leadership of key institutions in advancing AI-driven methodologies for SCRA.
@@ 4.10.5 Journals
Figure 13 illustrates the bibliographic coupling analysis of the core academic journals that are driving the advancements in SCRA research by showing the interconnections and thematic clustering among 17 influential journals. Journals such as the “International Journal of Production Research” and “Annals of Operations Research” play a crucial role, providing foundational studies and methodologies that inform AI-driven approaches to SCRA. These clusters reveal thematic trends and emerging focus areas and facilitate knowledge dissemination by connecting researchers across disciplines.
@@
Table 4: Publication Trends by Journal (2015-2024)
{tblr}
hline1-2,24 = -, Source Title & 2016 2017 2018 2019 2020 2021 2022 2023 2024 Total
ANNALS OF OPERATIONS RESEARCH           2 4 1   7
COMPUTERS INDUSTRIAL ENGINEERING       1   1 2 2   6
INTERNATIONAL JOURNAL OF PRODUCTION RESEARCH       1     4 1   6
SUSTAINABILITY       1 1 1 1 1   5
WIRELESS COMMUNICATIONS MOBILE COMPUTING           1 3     4
INDUSTRIAL MANAGEMENT DATA SYSTEMS     1     1 1   1 4
EXPERT SYSTEMS WITH APPLICATIONS 1           1 1   3
INTERNATIONAL JOURNAL OF LOGISTICS MANAGEMENT           1 2     3
INTERNATIONAL JOURNAL OF PRODUCTION ECONOMICS       1 1     1   3
NEURAL COMPUTING APPLICATIONS   1       1 1     3
COMPUTATIONAL INTELLIGENCE AND NEUROSCIENCE             2     2
ENTERPRISE INFORMATION SYSTEMS         2         2
FUTURE GENERATION COMPUTER SYSTEMS       1     1     2
INTERNATIONAL JOURNAL OF DISTRIBUTED SENSOR NETWORKS       1 1         2
INTERNATIONAL JOURNAL OF OPERATIONS PRODUCTION MANAGEMENT             1 1   2
JOURNAL OF INTELLIGENT FUZZY SYSTEMS           2       2
KYBERNETES               2   2
APPLIED SCIENCES-BASEL           1     1 2
OPERATIONS MANAGEMENT RESEARCH             2     2
ALEXANDRIA ENGINEERING JOURNAL             1     1
HELIYON                 1 1
APPLIED SCIENCE AND ENGINEERING JOURNAL FOR ADVANCED RESEARCH                 1 1
@@ 4.11 Journal Publication Trends
Table 4 presents a comprehensive breakdown of published papers in various source titles from 2015 to 2024. These source titles encompass a range of journals and publications contributing to the field of SCRA. The table clearly shows the publication trends, indicating how many papers were published in each source title each year and their cumulative total. It provides valuable insights into the distribution of research output and the source titles that have been active in this area.
@@ 5 Framework for SCRA using AI
Our proposed framework for AI-driven SCRA provides a comprehensive roadmap for SC professionals to integrate AI techniques effectively. Figure 14 visually shows the sequential steps involved, starting from defining key risk indicators to implementing preventive measures. By following these steps, practitioners can systematically define risk indicators, source and pre-process data, select and train models, and interpret results to facilitate informed decision-making. The framework is enriched with insights from existing studies on AI and ML applications in SC risk, presenting best practices, key considerations, and real-world challenges and opportunities identified in the literature.

### v5

@@ 4.2 Geographical Distribution of Research
Figure 3: Geographical distribution of the authors affiliated with 778 published papers on AI in SCRA topic.
Figure 3 visualizes the geographical distribution of research contributions. China leads with 39.49% of 778 publications, followed by the United States (11.72%) and England (11.59%). This dominance underscores the strong focus of Chinese institutions on AI-driven SCRA. The global spread of contributions emphasizes the universal relevance of SCRA, with increasing participation from emerging economies.
@@
Gupta S 14
Liu YK 12
Brintrup A 11
Kumar A 11
Mangla SK 8
Modgil S 8
Choi TM 7
Liu Y 7
Li J 6
Liu C 6
@@ 4.3 Analysis of Authors
As shown in Table 2, the top 10 authors have significantly influenced the domain of SCRA using AI/ML. Gupta S leads the field with 1.78% of the 778 publications, followed by Liu YK (1.53%) and Brintrup A (1.40%), who have contributed substantially to advancing AI-driven risk assessment. Other key contributors include Kumar A (1.40%), Mangla SK (1.02%), and Modgil S (1.02%), each demonstrating a strong research footprint in this domain. Additionally, Choi TM (0.89%), Liu Y (0.89%), Li J (0.76%), and Liu C (0.76%) round out the top 10 authors, further emphasizing the interdisciplinary and global nature of research in AI-based SCRA. These authors have played a crucial role in developing AI-driven risk assessment frameworks, predictive models, and decision-support systems, shaping the future of SC resilience and risk mitigation strategies.
@@ 4.4 Publisher Contributions
Figure 4: Publisher-wise distribution of 778 articles.
Figure 4 presents the contributions of various publishers. Elsevier is the most prominent, contributing over 262 articles, followed by Springer Nature, MDPI, Taylor & Francis, and IEEE. These publishers serve as key platforms for disseminating high-quality research, reflecting their alignment with the evolving needs of AI and SCRA scholarship.
@@ 4.5 Co-citation Analysis
Co-citation analysis was conducted to explore the intellectual structure of AI-SCRA research, focusing on authors, references, and publication sources. Figure 5(a) presents an author co-citation network comprising 55 authors grouped into three distinct clusters. The visualization highlights key contributors to AI-driven SCRA, with Ivanov and Baryannis playing central roles in the field. Other notable researchers, such as Dubey, Choi, and Zhu, also exhibit strong influence within their respective clusters. Figure 5(b) presents a co-citation network with three main clusters, highlighting key contributions to AI-driven SCRM. Notable references, including Ivanov (2019, 2020), Baryannis (2019), and Tang (2006), emphasize hybrid AI models, risk prediction frameworks, and post-COVID resilience strategies. Lastly, Figure 5(c) underscores the centrality of academic journals, with the International Journal of Production Research standing out as a key publication venue, reflecting its substantial impact on the field’s scholarly discourse and research progression.
@@ 4.6 Co-authorship Analysis
Co-authorship analysis is a powerful method for understanding collaborations among authors and researchers based on joint publication activities. This analysis highlights the collaborative nature of the research domain, uncovering influential partnerships and the significant contributions of specific countries. Figure 6(a) illustrates the co-authorship network among authors in this field. Using a threshold of at least one document and a minimum of 55 citations per author, 48 authors out of a total of 304 were identified as significant contributors. The visualization, generated using VOSviewer, reveals four distinct clusters, color-coded as red, green, blue, and yellow. These clusters represent groups of authors with strong collaborative ties. For example, the robust connections surrounding “Modgil” in the red cluster indicate frequent co-authorship and shared research initiatives, reflecting a high degree of collaboration within the cluster. Figure 6(b) showcases the co-authorship network at the country level. By applying thresholds of at least two documents and nine citations per country, 18 out of 36 countries were included in the analysis. The visualization highlights China, USA, and England as leading contributors, dominating document count, citation impact, and total link strength. This dominance underscores their pivotal roles in advancing AI/ML-driven SCRA research. The results also emphasize strong international collaborations between these countries and other key players, fostering a globally interconnected research ecosystem.
@@ 4.7 Co-occurrence Analysis
The co-occurrence analysis offers valuable insights into the thematic structure of SCRA and AI/ML research by identifying how keywords are interconnected within the scholarly literature. Using VOSviewer, we conducted this analysis by focusing on the units of analysis, such as “Authors’ Keywords” and “Keywords Plus.” This approach reveals latent connections and thematic clusters, enabling a deeper understanding of this domain’s critical topics and prevalent themes. The network map shown in Figure 7(a) illustrates the relationships among the 26 most frequently occurring author keywords. These keywords were selected from an initial pool of 1,483, applying a minimum threshold of three occurrences. Larger nodes, such as “artificial intelligence” and “machine learning,” indicate higher frequencies. The proximity of the nodes reflects their degree of association, with closely grouped nodes forming thematic clusters. Four clusters emerged, highlighting key themes. Figure 7(b) illustrates the co-occurrence of keywords in the reviewed literature. The analysis identified 47 frequently occurring terms from 1,032, using a minimum threshold of three occurrences. These keywords were categorized into five distinct clusters, reflecting prevalent themes and connections within the research field. Terms such as “management” and “framework” exhibited the strongest linkages, indicating their crucial roles in connecting different thematic areas. Additionally, other significant clusters highlighted emerging trends, including “supply chain resilience”, “big data”, “blockchain” and “sustainability”. This co-occurrence analysis offers a comprehensive overview of the intellectual structure of SCRA research, showcasing established and emerging themes. The findings emphasize the interconnectivity of concepts and provide guidance for future research directions by identifying critical areas of scholarly focus.
@@
Affiliations Published paper count % of 778 articles
HONG KONG POLYTECHNIC UNIVERSITY 22 2.799
CHINESE ACADEMY OF SCIENCES 21 2.672
INDIAN INSTITUTE OF TECHNOLOGY (IIT) SYSTEM 20 2.545
UNIVERSITY OF TEHRAN 19 2.417
INDIAN INSTITUTE OF MANAGEMENT (IIM) SYSTEM 15 1.908
NEOMA BUSINESS SCH 15 1.908
UNIVERSITY OF CAMBRIDGE 15 1.908
HEBEI UNIVERSITY 14 1.781
EMLYON BUSINESS SCHOOL 11 1.399
INDIAN INSTITUTE OF TECHNOLOGY (IIT) DELHI 11 1.399
Others 615 79.26
@@ 4.8 Top 10 Affiliations
Table 3 highlights the top 10 affiliations contributing to SCRA research, ranked by their published paper count. Hong Kong Polytechnic University leads the list with 22 publications (2.80%), followed closely by the Chinese Academy of Sciences (2.67%) and the Indian Institute of Technology (IIT) System (2.55%). Other key contributors include the University of Tehran (2.42%), the Indian Institute of Management (IIM) System (1.91%), and NEOMA Business School (1.91%), all of which have played a crucial role in advancing AI-driven SCRA methodologies. Prominent European institutions such as the University of Cambridge (1.91%), Hebei University (1.78%), and Emlyon Business School (1.40%) further highlight the global nature of SCRA research, emphasizing interdisciplinary contributions across Asia, Europe, and the Middle East. Notably, IIT Delhi (1.40%) is a significant research hub in this field. Together, these top 10 affiliations account for a substantial share of high-impact studies in SCRA, driving advancements in AI applications for risk mitigation. Meanwhile, the “Others” category, comprising 615 institutions, represents 79.26% of the contributions, reflecting the diverse and widespread nature of research in this domain.
@@ 4.9.1 Authors
Figure 8 illustrates the bibliographic coupling analysis of 28 authors, grouping them into five thematic clusters based on shared references, highlighting intellectual connections in SCR research. Prominent contributors like Lim M.K., Gupta, Shivam, Wang G.-J., and Hussain F.K. demonstrate centrality, reflecting their significant influence in areas such as risk prediction, disruption management, and quantitative modeling. The clusters reveal interdisciplinary themes, including resilience strategies, industry-specific applications, and the integration of AI-driven methods. This analysis showcases established research directions and emerging trends, providing a roadmap for future studies and collaborations in SCR.
@@ 4.9.2 Countries
Figure 9 illustrates the bibliographic coupling among 18 countries, categorizing them into clusters highlighting global research collaborations in AI-driven SCRA. Countries such as China, the United States, and England emerge as central hubs, showcasing their pivotal roles in advancing research and promoting international partnerships. The clusters signify thematic alignments across regions, with nations collaborating on critical topics like predictive analytics, sustainability, and resilience strategies.
@@ 4.9.3 Documents
Figure 10 illustrates the bibliographic coupling analysis of 62 documents, visually depicting the thematic clustering and citation relationships among the studies. The eight clusters highlight distinct research themes within the field, showcasing the interconnectedness of key works. Highly influential documents, such as those by “Baryannis G.”, “Cavalcante I. M.”, and “Zhu Q.”, emerge as central nodes, reflecting their significant contributions to the discourse on SCRA using AI. This visualization underscores the collaborative and evolving nature of the research domain, providing insights into how various studies align and contribute to specific thematic advancements.
@@ 4.9.4 Organizations
Figure 11 presents the bibliographic coupling analysis of 24 organizations, highlighting their collaborative and thematic alignments in AI-based SCRA research. The visualization categorizes institutions into three clusters, highlighting key contributors such as the Hong Kong Polytechnic University, the University of Cambridge, and Nanjing University of Posts and Telecommunications. This analysis underscores the global distribution of research efforts and the leadership roles of prominent institutions in advancing AI-driven methodologies for SCRA.
@@ 4.9.5 Journals
Figure 12 illustrates the bibliographic coupling analysis of the core academic journals driving the advancements in SCRA research by showing the interconnections and thematic clustering among 17 influential journals. Journals such as the “International Journal of Production Research” and “Annals of Operations Research” play a crucial role, providing foundational studies and methodologies that inform AI-driven approaches to SCRA. These clusters reveal thematic trends and emerging focus areas and facilitate knowledge dissemination by connecting researchers across disciplines.
@@
Table 4: Publication Trends by Journal (2015-2025)
Source Title Publication Year Grand Total
2015 2016 2017 2018 2019 2020 2021 2022 2023 2024 2025
INTERNATIONAL JOURNAL OF PRODUCTION RESEARCH 1 5 3 4 18 7 38
SUSTAINABILITY 1 2 4 5 9 9 3 1 34
ANNALS OF OPERATIONS RESEARCH 2 10 18 1 31
ENGINEERING APPLICATIONS OF ARTIFICIAL INTELLIGENCE 1 1 2 2 5 8 2 21
INTERNATIONAL JOURNAL OF PRODUCTION ECONOMICS 1 1 1 1 1 6 8 2 21
COMPUTERS & INDUSTRIAL ENGINEERING 2 1 3 2 3 5 4 20
IEEE ACCESS 2 4 1 2 4 3 16
IEEE TRANSACTIONS ON ENGINEERING MANAGEMENT 4 12 16
EXPERT SYSTEMS WITH APPLICATIONS 1 1 1 2 3 4 3 15
TRANSPORTATION RESEARCH PART E-LOGISTICS AND TRANSPORTATION REVIEW 1 1 1 3 1 4 11
JOURNAL OF CLEANER PRODUCTION 2 5 1 2 10
APPLIED SCIENCES-BASEL 1 2 1 5 9
INDUSTRIAL MANAGEMENT & DATA SYSTEMS 1 3 1 2 7
JOURNAL OF INTELLIGENT & FUZZY SYSTEMS 1 2 3 1 7
PRODUCTION AND OPERATIONS MANAGEMENT 4 1 2 7
TECHNOLOGICAL FORECASTING AND SOCIAL CHANGE 1 2 2 2 7
APPLIED SOFT COMPUTING 1 1 1 1 2 6
DECISION SUPPORT SYSTEMS 1 3 2 6
EUROPEAN JOURNAL OF OPERATIONAL RESEARCH 1 1 2 1 1 6
INDUSTRIAL MARKETING MANAGEMENT 1 1 1 3 6
INTERNATIONAL JOURNAL OF LOGISTICS-RESEARCH AND APPLICATIONS 3 3 6
INTERNATIONAL JOURNAL OF PHYSICAL DISTRIBUTION & LOGISTICS MANAGEMENT 1 2 1 2 6
OPERATIONS MANAGEMENT RESEARCH 3 1 2 6
PLOS ONE 1 1 1 2 1 6
WIRELESS COMMUNICATIONS & MOBILE COMPUTING 3 3 6
Others 2 2 4 9 27 28 46 77 68 151 40 454
Grand Total 4 5 5 13 39 48 80 124 132 256 72 778
@@ 4.10 Journal Publication Trends
Table 4 presents a comprehensive breakdown of published papers in various source titles from 2015 to 2025. These source titles encompass a range of journals and publications contributing to the field of SCRA. The table clearly shows the publication trends, indicating how many papers were published in each source title each year and their cumulative total. It provides valuable insights into the distribution of research output and the source titles that have been active in this area.
@@ 5 Framework for SCRA using AI
Our proposed framework for AI-driven SCRA provides a comprehensive pipeline for SC professionals to integrate AI techniques effectively. Figure 13 visually shows the sequential steps involved, from defining key risk indicators to implementing preventive measures. Following these steps, practitioners can systematically define risk indicators, source and pre-process data, select and train models, and interpret results to facilitate informed decision-making. The framework is enriched with insights from existing studies on AI and ML applications in SC risk, presenting best practices, key considerations, and real-world challenges and opportunities identified in the literature.